OpenStudy (taylormarie213):

70% of what number is 224?

OpenStudy (anonymous):

\[\frac{ 70 }{ 100 } = \frac{ 224 }{ x }\]

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

id we take that number to be x, it can be written in mathematical terms to be: \[\frac{ 70 }{ 100 }\times x = 224\] to find x, \[x=\frac{ 224 \times 100 }{ 70 }\] \[x=??\]

OpenStudy (anonymous):

if..

OpenStudy (anonymous):

70x= 224(100)

OpenStudy (anonymous):

70x= 22400

OpenStudy (anonymous):

divide both sides by 70

OpenStudy (anonymous):

x= 320